Aggregate Expenditure Question

Assume the following, m1 = .15, i1 = .15, and c1 = .8. The multiplier in this case is:

a) 5

b) .8

c) 4

d) 2.86
Answer #1

Multiplier = 1 / ( 1 - c + i - m )

= 1 / ( 1 - .8 + .15 - .15 ) = 1 / ( .2) = 5

Ans. A. Multiplier = 5

m1 is propensity to import

c1 is propensity to consume

i1 is propensity to invest

Multiplier is change in output due to change in autonomous variables.